# Harnessing Knowledge Retrieval with Large Language Models for Clinical Report Error Correction

### v1

@@ Abstract
This study proposes an approach for error correction in clinical radiology reports, leveraging large language models (LLMs) and retrieval-augmented generation (RAG) techniques. The proposed framework employs internal and external retrieval mechanisms to extract relevant medical entities and relations from the report and external knowledge sources. A three-stage inference process is introduced, decomposing the task into error detection, localization, and correction subtasks, which enhances the explainability and performance of the system. The effectiveness of the approach is evaluated using a benchmark dataset created by corrupting real-world radiology reports with realistic errors, guided by domain experts. Experimental results demonstrate the benefits of the proposed methods, with the combination of internal and external retrieval significantly improving the accuracy of error detection, localization, and correction across various state-of-the-art LLMs. The findings contribute to the development of more robust and reliable error correction systems for clinical documentation.
Harnessing Knowledge Retrieval with Large Language Models for Clinical Report Error Correction
Jinge Wu1thanks: These authors contributed equally to this work., Zhaolong Wu211footnotemark: 1, Abul Hasan1, Yunsoo Kim1, Jason P.Y. Cheung2thanks: Corresponding authors., Teng Zhang [^2] footnotemark: 2, Honghan Wu1322footnotemark: 2 1University College London 2University of Hong Kong 3University of Glasgow {wuzl01}@connect.hku.hk, {cheungjp, tgzhang}@hku.hk, {jinge.wu.20, honghan.wu}@ucl.ac.uk
@@ 1 Introduction
To rigorously evaluate the performance of our proposed approach, we establish a benchmark for the clinical report error detection task by intentionally introducing common mistakes into real-world electronic health records, with the guidance of domain experts. This benchmark enables a realistic assessment of the error detection capabilities of our system and provides a standardized testing ground for future research in this area.
Through extensive experiments, we conduct a thorough assessment of the effectiveness of our proposed approaches, demonstrating their performance through comprehensive evaluations and analyses. By comparing our system to state-of-the-art baselines and ablating various components of our framework, we aim to provide a clear understanding of the strengths and limitations of our approach, as well as identify potential avenues for future improvement.
@@
By harnessing the power of LLMs while addressing their limitations through a decomposed, multi-stage approach and reasoning mechanisms, our system aims to enhance the accuracy and reliability of clinical documentation, ultimately contributing to improved healthcare delivery and patient outcomes.
@@ 3.1 Internal and External Retrieval
Our work focuses on radiology report error correction which contains separate sections for Findings and Impression. The Findings section details observations from a radiograph, while the Impression section provides diagnostic conclusions drawn from those findings. Inspired by Yan et al. (2023), we transform this report into a graph representation using RadGraph. In the resulting graph, nodes correspond to two main types of entities: anatomical entities (e.g., lungs, ribs) and observational entities (e.g., clear, acute). Edges connecting these nodes denote relations such as modification (an observational entity modifying an anatomical entity), location (specifying the anatomical location of an observation), or suggestion (an observation suggesting a particular diagnosis).
The extraction of these key medical mentions and their relationships is expected to enable more targeted reasoning by the model, as it can focus on the most salient information while considering the connections between various entities. Moreover, RadGraph provides additional contextual information by categorizing observations based on their presence or absence in the report. Observations are denoted as either "Observation: Definitely Present" (OBS-DP) or "Observation: Definitely Absent" (OBS-DA). This explicit labeling of the presence or absence of observations enhances the input for the model, allowing it to better identify contextual inconsistencies within the report. By leveraging this rich graph-based representation of the clinical report, our method aims to improve the accuracy and consistency of the model’s reasoning process.
@@ 3.2 Rephrasal Augmentation
When applying RadGraph to extract entities and relations from radiology reports, we further rephrase the extracted information into concise sentences to enhance interpretability and downstream reasoning capabilities. RadGraph proposes a principled relation extraction schema to reconstruct radiology report findings from extracted clinical entities and their semantic relations. The schema defines three relation types: suggestive_of, located_at, and modify. These can be used to rephrase the entities and relations into sentences. For instance, consider a scenario where we have two relation triples: $\langle lower,modify,lobe\rangle$ and $\langle opacity,located\_at,lobe\rangle$. The first triple indicates that the entity "lower" modifies the entity "lobe," while the second triple signifies that the entity "opacity" is located at the entity "lobe." By combining these two triples based on their common entity "lobe," we can generate a meaningful phrase: "lower lobe opacity." This phrase captures the essential information conveyed by the individual relation triples, presenting it in a more concise and understandable manner. Furthermore, our method takes into account the negation of mentions, which is crucial for accurately representing the absence of certain findings. In cases where a mention is negated, we incorporate the word "no" at the beginning of the generated phrase. For example, if the relation triple $\langle opacity,located\_at,lobe\rangle$ is negated, we would generate the phrase "no lobe opacity."
@@
3.3 Three-Stage Inference
To enhance the inference capabilities and explainability of error correction in clinical reports, similar to Ben Abacha et al. (2024), we propose the three-stage approach by decomposing the task into three distinct subtasks: error detection, error localization, and error correction. By addressing these subtasks separately, the model can focus on smaller, more manageable objectives, leading to improved performance and increased interpretability.
In the error detection stage, the model aims to identify whether an error is present in the given clinical text. This binary classification task enables the model to distinguish between correct and incorrect information by learning the characteristics that indicate the presence of errors, such as inconsistencies, contradictions, or deviations from standard medical terminology. If an error is detected, the model proceeds to the error localization stage, where it pinpoints the precise location of the error within the text. This stage involves a fine-grained analysis of the clinical report, requiring the model to identify the specific words, phrases, or sentences that contain the error. By accurately determining the scope and nature of the error, the model can provide targeted corrections. In the final error correction stage, the model generates the corrected sentence based on the information gathered from the previous stages. By leveraging the knowledge of the error’s presence and location, the model can create a contextually appropriate correction that maintains the overall coherence and accuracy of the clinical report.
The three-stage approach offers several advantages over the end-to-end error correction task. By decomposing the task into smaller, focused subtasks, the model can learn more efficiently and effectively, capturing the intricacies of error detection and correction more accurately. This process allows the model to respond with higher accuracy by focusing on each subtask independently before combining the results. Furthermore, this approach helps to reduce hallucination by encouraging the model to think through the task step by step, mitigating the risk of generating irrelevant or inconsistent corrections. Moreover, the three-stage approach enhances the explainability of the model. By separating the error detection, localization, and correction processes, it becomes easier to interpret the model’s decisions and understand the reasoning behind its predictions. Each stage provides clear insights into the model’s thought process, from identifying the presence of an error to pinpointing its location and generating a suitable correction.
@@ 4.1 Dataset
Our study utilizes chest radiology reports from the MIMIC-CXR dataset Johnson et al. (2019) as an exemplar. MIMIC-CXR is a large-scale collection of real-world radiology data, providing a valuable resource for our analysis. From this dataset, we employ 112,251 radiology reports as reference data, representing correctly annotated ground truth reports. Additionally, we curate 1,622 reports as evaluation data, where a subset of these reports are purposefully corrupted to simulate real-world annotation errors (see details in Table 1).
Reference data will be used for external retrieval to provide augmented knowledge and clinical findings for error detection and correction. Evaluation data, on the other hand, refers to the data used for evaluation of the task, where some of the reports contain errors.
To introduce realistic errors into the evaluation data, we employ two strategies with the guidance of clinical experts and radiologists. Firstly, we replace essential clinical observations with irrelevant ones. Specifically, we utilize the 12 important observations from the CheXpert dataset Irvin et al. (2019) (excluding "No Finding" and "Support Devices" from the original 14 observations) and substitute them with unrelated observations carefully selected to be clinically irrelevant to the report context. Secondly, we selectively alter the polarity of observations by removing negation words such as "no" or "no evidence of", effectively transforming negated findings into affirmed ones. This process aims to mimic common annotation mistakes where the presence or absence of clinical findings is incorrectly recorded. By corrupting a subset of the evaluation data through these two strategies, we aim to create a more challenging and realistic benchmark that captures the types of errors that may arise in practical radiology report annotation scenarios.
@@ 5 Result
Three-stage Inference End-to-End
Error Detection Error Localization Error Correction Error Correction
ACC ACC AggNLG AggNLG
MMedLM2 41.49 30.94 58.27 (+10.47) 47.80
Aloe 45.31 43.34 90.17 (+26.84) 63.33
Phi3-mini 67.26 47.71 74.08 (-0.28) 74.36
Phi3-small 79.03 63.44 86.57 (+6.53) 80.03
Phi3-medium 73.67 69.73 90.25 (+5.79) 84.47
Llama3-8B 37.79 37.29 94.29 (+9.95) 84.34
@@
5.1 Comparisons with Three-Stage Inference
Table 2 presents a comparison of the performance of various language models on the tasks of error detection, localization, and correction using three-stage inference and end-to-end approaches. For the three-stage inference approach, the accuracy scores for error detection range from 37.79% (Llama3-8B) to 79.03% (Phi3-small). Error localization accuracy follows a similar trend, with Llama3-8B achieving the lowest score (37.29%) and Phi3-medium obtaining the highest (69.73%). In terms of error correction, the aggregate NLG score (AggNLG) is used to assess performance, where AggNLG is the averagte score of ROUGE-1, BERTScore, and BLEURT. Llama3-8B achieves the highest AggNLG score of 94.29%, followed closely by Phi3-medium and Aloe. The end-to-end error correction approach, yields similar pattern compared with the three-stage inference approach, where Llama3-8B and Phi3-medium achieves the top two AggNLG scores.
Comparing the three-stage inference and end-to-end approaches, it is evident that the three-stage inference generally yields higher AggNLG scores for error correction. The two medical LLMs, MMedLM2 and Aloe, improves the most with 10.47% and 26.84% respectively. Phi3-mini does not show improvements in this case, with a slight decrease of 0.28%. The overall performance suggests that breaking down the error correction process into distinct stages of detection, localization, and correction can lead to better overall performance compared to an end-to-end approach.
Another point worth noticing is that the calculation of error correction requires the model to analyze the input radiological report: if an error is detected, the model will correct it by generating the amended report entirety; otherwise, it will return "NA". The scoring process is conducted exclusively when both the model and the ground truth identify errors (neither being "NA"), and the rectified results are compared. Radiological report errors typically involve minor corrections of a few words, leaving the majority of the report content unaltered, which explains the generally higher NLG scores in the result Tables.
@@ 5.2 Comparisons with Internal and External Retrieval
Table 3: Comparisons of internal and external retrieval on error detection (accuracy).
In Table 3, which focuses on error detection accuracy, where all models significantly improves while applying both internal and external retrieval. MMedLM2 improves the most by 31.56% and Phi3-small improves the least by 1.18%. Additionally, it is observed that using only internal retrieval does not improve the performance significantly. Only Phi3-mini and Llama3-8B shows improvement with internal retrieval.
@@
Table 4: Comparisons of internal and external retrieval on error localization (accuracy).
Table 4 presents results for error localization accuracy. All models show improvement while using both internal and external retrieval, except for Phi3-medium. Llama3-8B and MMedLM2 improve the most by more than 15%. Similar to error detection, it is noticed that internal retrieval does not provide effectiveness for this task as well, where only Phi3-small shows increased accuracy with internal retrieval.
@@
Overall, these tables demonstrate the effectiveness of internal and external retrieval methods in enhancing the performance of language models on error detection, localization, and correction tasks. The results suggest that incorporating retrieval mechanisms can lead to significant improvements in accuracy and AggNLG scores compared to the baseline models.
@@ 6 Discussion
In conclusion, the combined retrieval method proves to be a valuable approach for enhancing the performance of language models on error detection, localization, and correction tasks. However, the effectiveness of this method varies depending on the underlying model architecture, training data, and the clarity of instructions provided. While the Llama3-8B model excels at following instructions, the Phi3 models demonstrate strong performance in zero-shot scenarios. Fine-tuning models on domain-specific data, such as medical text, may impact their language capabilities, and careful consideration should be given to the training data and prompts used during the fine-tuning process. Future research should explore methods to improve the models’ understanding and utilization of summary information, as well as investigate strategies to balance the benefits of domain-specific fine-tuning with the preservation of general language capabilities.
7 Conclusion
In conclusion, this study introduces a novel framework for error detection, localization, and correction in clinical radiology reports, combining the strengths of LLMs and RAG techniques. The proposed approach leverages internal and external retrieval mechanisms to extract relevant medical entities and relations, enhancing the contextual understanding of the LLMs. By decomposing the task into three distinct subtasks through a three-stage inference process, the system achieves improved performance and explainability compared to end-to-end approaches. Experimental results on a benchmark dataset, created by corrupting real-world radiology reports with realistic errors, demonstrate the effectiveness of the proposed methods. The combination of internal and external retrieval significantly improves the accuracy of error detection, localization, and correction across various state-of-the-art LLMs.

### v2

# Integrating Knowledge Retrieval and Large Language Models for Clinical Report Correction

@@ Abstract
This study proposes an approach for error correction in radiology reports, leveraging large language models (LLMs) and retrieval-augmented generation (RAG) techniques. The proposed framework employs a novel internal+external retrieval mechanism to extract relevant medical entities and relations from the report of interest and an external knowledge source. A three-stage inference process is introduced, decomposing the task into error detection, localization, and correction subtasks, which enhances the explainability and performance of the system. The effectiveness of the approach is evaluated using a benchmark dataset created by corrupting real-world radiology reports with realistic errors, guided by domain experts. Experimental results demonstrate the benefits of the proposed methods, with the combination of internal and external retrieval significantly improving the accuracy of error detection, localization, and correction across various state-of-the-art LLMs. The findings contribute to the development of more robust and reliable error correction systems for clinical documentation.
Integrating Knowledge Retrieval and Large Language Models for Clinical Report Correction
Jinge Wu1thanks: These authors contributed equally to this work., Zhaolong Wu211footnotemark: 1, Ruizhe Li3, Abul Hasan1, Yunsoo Kim1, Jason P.Y. Cheung2thanks: Corresponding authors., Teng Zhang [^2] footnotemark: 2, Honghan Wu1422footnotemark: 2 1University College London 2University of Hong Kong 3University of Aberdeen 4University of Glasgow {jinge.wu.20, honghan.wu}@ucl.ac.uk
@@ 3 Method
3.1 Three-Stage Inference
To enhance the inference capabilities and explainability of error correction in clinical reports, we propose a three-stage approach inspired by Ben Abacha et al. (2024). This method decomposes the task into three distinct subtasks, allowing the model to focus on smaller, more manageable objectives, leading to improved performance and increased interpretability.
Our three-stage approach consists of:
  • Error Detection: The model identifies whether an error is present in the given clinical text. This binary classification task enables the model to distinguish between correct and incorrect information by learning the characteristics that indicate the presence of errors.
  • Error Localization: If an error is detected, the model pinpoints its precise location within the text. This stage involves a fine-grained analysis of the clinical report to identify the specific words, phrases, or sentences that contain the error.
  • Error Correction: Based on the information gathered from the previous stages, the model generates the corrected sentence. This stage leverages the knowledge of the error’s presence and location to create a contextually appropriate correction that maintains the overall coherence and accuracy of the clinical report.
This approach offers several advantages over end-to-end error correction. By decomposing the task, the model can learn more efficiently and effectively, capturing the intricacies of error detection and correction more accurately. It also enhances explainability by providing clear insights into the model’s thought process at each stage. Furthermore, this step-by-step method helps reduce hallucination and encourages more targeted, accurate corrections.
@@ 3.2 Internal and External Retrieval
Here we introduce a novel graph and knowledge-based retrieval method for prompt engineering to enhance our system’s reasoning capabilities across all phases. This approach simplifies complex information and eliminates redundancies, facilitating more efficient and effective reasoning. Our study focuses on error correction in radiology reports, specifically addressing the Findings and Impression sections. The Findings section documents radiographic observations, while the Impression section provides diagnostic conclusions based on these findings. Our retrieval technique incorporates both internal and external retrieval processes to enhance the accuracy and consistency of radiology report error correction. The internal retrieval transforms the report into a graph representation using RadGraph, focusing on key medical entities and their relationships, while the external retrieval leverages a large-scale dataset of clinical reports to provide supplementary knowledge and context for more robust error detection.
@@
Inspired by Yan et al. (2023), we transform the report into a graph representation using RadGraph. In the resulting graph, nodes correspond to two main types of entities: anatomical entities (e.g., lungs, ribs) and observational entities (e.g., clear, acute). Edges connecting these nodes denote relations such as modification (an observational entity modifying an anatomical entity), location (specifying the anatomical location of an observation), or suggestion (an observation suggesting a particular diagnosis).
The extraction of these key medical mentions and their relationships is expected to enable more targeted reasoning by the model, as it can focus on the most salient information while considering the connections between various entities. Moreover, RadGraph provides additional contextual information by categorizing observations based on their presence or absence in the report. Observations are denoted as either “Observation: Definitely Present" (OBS-DP) or “Observation: Definitely Absent" (OBS-DA). This explicit labeling of the presence or absence of observations enhances the input for the model, allowing it to better identify contextual inconsistencies within the report. By leveraging this rich graph-based representation of the clinical report, our method aims to improve the accuracy and consistency of the model’s reasoning process.
@@ 3.3 Rephrasal Augmentation
When applying RadGraph to extract entities and relations from radiology reports, we further rephrase the extracted information into concise sentences to enhance interpretability and downstream reasoning capabilities. RadGraph proposes a principled relation extraction schema to reconstruct radiology report findings from extracted clinical entities and their semantic relations. The schema defines three relation types: suggestive_of, located_at, and modify. These can be used to rephrase the entities and relations into sentences. For instance, consider a scenario where we have two relation triples: $\langle lower,modify,lobe\rangle$ and $\langle opacity,located\_at,lobe\rangle$. The first triple indicates that the entity lower modifies the entity lobe, while the second triple signifies that the entity opacity is located at the entity lobe. By combining these two triples based on their common entity lobe, we can generate a meaningful phrase: lower lobe opacity. This phrase captures the essential information conveyed by the individual relation triples, presenting it in a more concise and understandable manner. Furthermore, our method takes into account the negation of mentions, which is crucial for accurately representing the absence of certain findings. In cases where a mention is negated, we incorporate the word no at the beginning of the generated phrase. For example, if the relation triple $\langle opacity,located\_at,lobe\rangle$ is negated, we would generate the phrase no lobe opacity.
@@ 4.1 Dataset
Our study utilizes real-world clinical reports from the MIMIC-CXR dataset Johnson et al. (2019). MIMIC-CXR is a large-scale collection of ICU hospital data, providing a valuable resource for our analysis. From this dataset, we employ 112,251 radiology reports as reference data, representing correctly annotated ground truth reports. Additionally, we curate 1,622 radiology reports as evaluation data, where a subset of these reports are purposefully corrupted to simulate real-world annotation errors (see details in Table 1).
Reference data will be used for external retrieval to provide augmented knowledge and clinical findings for error detection and correction. Evaluation data, on the other hand, refers to the data used for evaluation of the task, where some of the reports contain errors. More data statistics can be found in Appendix (Table 10)
To introduce realistic errors into the evaluation data, we employ two strategies with the guidance of clinical experts and radiologists. Firstly, we replace essential clinical observations with irrelevant ones. Specifically, we utilize the 12 important observations from the CheXpert dataset Irvin et al. (2019) (excluding “No Finding" and “Support Devices" from the original 14 observations) and substitute them with unrelated observations carefully selected to be clinically irrelevant to the report context (see the lists in Table 8 and 9 in Appendix). Secondly, we selectively alter the polarity of observations by removing negation words such as “no" or “no evidence of", effectively transforming negated findings into affirmed ones. This process aims to mimic common annotation mistakes where the presence or absence of clinical findings is incorrectly recorded. By corrupting a subset of the evaluation data through these two strategies, we aim to create a more challenging and realistic benchmark that captures the types of errors that may arise in practical radiology report annotation scenarios.
@@ 4.2 Evaluation
The calculation of error correction requires the model to analyze the input radiological report: if an error is detected, the model will correct it by generating the amended report entirety; otherwise, it will return “NA". The scoring process is conducted exclusively when both the model and the ground truth identify errors (neither being “NA"), and the rectified results are compared. Radiological report errors typically involve minor corrections of a few words, leaving the majority of the report content unaltered, which explains the generally higher NLG scores in the result tables.
@@ 5 Result
5.1 Comparisons with Three-Stage Inference
Table 2 presents a comparison of the performance of various language models on the tasks of error correction using three-stage inference and end-to-end approaches. The results demonstrate a clear trend: the three-stage inference method generally yields superior AggNLG scores compared to end-to-end approaches across most models examined.
MMedLM2 and Aloe, both medical language models, show the most substantial gains of 10.47% and 26.84% respectively. Interestingly, Phi3-mini exhibits a slight performance decrease (-0.28%), suggesting that the efficacy of this approach may depend on model characteristics. Larger size models, including Phi3 variants (small and medium) and Llama3-8B, demonstrate moderate to significant improvements ranging from 4.53% to 9.95%. This trend indicates a possible positive correlation between model scale and the effectiveness of the three-stage approach, potentially due to larger models’ enhanced ability to leverage task decomposition.
Overall, these findings suggest that decomposing the error correction process into distinct stages of detection, localization, and correction can lead to enhanced performance compared to end-to-end approaches. This improvement is likely attributable to the structured nature of the three-stage method, which allows models to focus on specific aspects of the task sequentially, potentially mitigating error propagation and enabling more precise corrections.
End-to-End Three-stage Inference
MMedLM2 47.80 58.27 (+10.47)
Aloe 63.33 90.17 (+26.84)
Phi3-mini 74.36 74.08 (-0.28)
Phi3-small 80.03 86.57 (+6.53)
Phi3-medium 84.47 90.25 (+5.79)
Llama3-8B 84.34 94.29 (+9.95)
@@ 5.2 Comparisons with Internal and External Retrieval
Table 3: Comparisons of internal and external retrieval on error detection (acc%).
Table 3 shows the accuracy on error detection. All models demonstrate significant improvements when employing both internal and external retrieval. MMedLM2 exhibits the most substantial enhancement, with a 31.56% increase, while Phi3-small shows the least improvement at 1.18%. Notably, internal retrieval alone yields limited benefits, with only Phi3-mini and Llama3-8B showing marginal improvements.
@@
Table 4: Comparisons of internal and external retrieval on error localization (acc%).
Table 4 illustrates error localization accuracy. All models show improvement while using both internal and external retrieval, except for Phi3-medium. Llama3-8B and MMedLM2 improve the most by more than 15%. Similar to error detection, it is noticed that internal retrieval does not provide effectiveness for this task as well, where only Phi3-small shows increased accuracy with internal retrieval.
@@
Error Detection (acc %) Error Localization (acc %) Error Correction (AggNLG)
Simple RAG Ours Simple RAG Ours Simple RAG Ours
MMedLM2 52.65 73.05 (+20.4) 35.88 46.05 (+10.17) 48.13 53.50 (+5.37)
Aloe 58.81 67.26 (+8.45) 44.51 51.35 (+6.84) 74.28 74.77 (+0.49)
Phi3-mini 65.59 73.06 (+7.47) 42.97 52.65 (+9.68) 75.76 78.85 (+3.09)
Phi3-small 72.51 80.21 (+7.7) 58.93 65.04 (+6.11) 84.56 86.67 (+2.11)
Phi3-medium 73.13 79.04 (+5.91) 60.91 63.44 (+2.53) 88.87 92.25 (+3.38)
Llama3-8B 51.84 62.27 (+10.43) 46.92 53.14 (+6.22) 88.34 94.43 (+6.09)
Table 6: Comparisons of our method with simple RAG. Our methods refer to the internal+external retrieval.
Simple RAG Ours
MMedLM2 30.51 22.0 -8.51
Aloe 25.6 16.1 -9.55
Phi3-mini 24.1 15.5 -8.58
Phi3-small 25.0 15.9 -9.11
Phi3-medium 31.4 21.9 -9.45
LLAMA3 8B 23.2 14.5 -8.68
Table 7: Inference efficiency comparison with simple RAG and our method (seconds taken per instance).
5.3 Comparison with Simple RAG
To further evaluate the effectiveness of our proposed method, we conducted a comparative analysis against a simple Retrieval-Augmented Generation (RAG) approach. While simple RAG typically involves retrieving relevant documents from a reference corpus and providing them as additional context to the language model, our method combines both internal and external retrieval mechanisms, focusing on key medical entities and their relationships.
Table 6 presents a comparison of our method with simple RAG across the tasks of error detection, localization, and correction. The results demonstrate that our method consistently outperforms simple RAG across all evaluation metrics. In error detection, our approach shows an average improvement of 10.06 percentage points over simple RAG. Notably, the MMedLM2 model exhibits the most significant enhancement, with an increase of 20.4 percentage points. For the error localization task, our method achieves an average improvement of 6.92 percentage points, with the Phi3-mini model showing the most substantial gain of 9.68 percentage points. In terms of error correction, our method consistently yields improvements in AggNLG scores. Although the magnitude of improvement is relatively smaller compared to the detection and localization tasks, all models demonstrate positive changes. The Llama3-8B model shows the most significant improvement, with an increase of 6.09 points in AggNLG score.
Table 7 compares the inference efficiency of our method against simple RAG. The results indicate that our method significantly outperforms simple RAG in terms of processing time per instance, with an average reduction of 9.15 seconds. This efficiency gain can be attributed to the more targeted information retrieval strategy in our method, enabling the model to locate and process relevant information more expeditiously.
These results underscore the advantages of our proposed retrieval method. By integrating internal and external retrieval mechanisms, our approach not only enhances the accuracy of error detection, localization, and correction but also significantly improves inference efficiency. This demonstrates that a well-designed retrieval strategy can lead to substantial performance improvements in the task of clinical report error correction.
The superior performance of our method can be attributed to several factors. Firstly, the graph-based internal retrieval allows for a more nuanced understanding of the report’s content, capturing intricate relationships between medical entities that might be overlooked in a simple text-based retrieval. Secondly, the external retrieval provides a broader medical context, potentially helping the model to identify subtle inconsistencies or errors that may not be apparent from the report alone.
@@ 6 Discussion
Moreover, we also conduct human evaluation for further analysis. More detail can be found in the Appendix. Our human evaluation results (Figure 3) reveal that the three-stage inference (internal + external) method significantly outperforms other approaches across accuracy, factual consistency, and clinical relevance. This improvement is particularly notable for models like Phi3-medium and Llama3-8B, with MMedLM2 showing dramatic increases in factual consistency and clinical relevance. These findings underscore the importance of human assessment in evaluating medical report generation models and demonstrate the enhanced capabilities of our proposed method across various model architectures.
In conclusion, our three-stage inference process, integrated with external knowledge retrieval, proves to be a robust strategy for improving accuracy, factual consistency, and clinical relevance in medical text analysis. This approach significantly improves models’ reasoning capabilities and task-specific performance without domain-specific fine-tuning, thereby enhancing efficiency and explainability while preserving general language capabilities.
Future research should focus on refining retrieval mechanisms, optimizing external knowledge integration, and enhancing models’ utilization of summary information. These advancements will be crucial in further improving the reliability and versatility of language models in healthcare applications.